OpenStudy (anonymous):

Give the maxium value of x if 12-3(x+1) is greater then or equal to 1/2(3-5)

OpenStudy (bahrom7893):

GO FOR IT MATT!

OpenStudy (bahrom7893):

I'm zoned out

OpenStudy (anonymous):

12-3x-3>=3/2-5/2 9-3x>=-2/2 9-3x>=-1 -3x>=-10 x<=10/3 maximum value 10/3

OpenStudy (anonymous):

i'm also zoned out aha. i keep coming back to open study as a break from my own not fun physics homework

OpenStudy (anonymous):

i'm also zoned out aha. i keep coming back to open study as a break from my own not fun physics homework